A company makes and sells a single product. The company employs a flexible budgeting system that covers a relevant range from 25,000 units to 30,000 units. Budget data for April, based on 27,000 units are as follows:

Prime costs $35 per unit
Manufacturing overhead $137,000 plus $4 per unit
Selling and admin expenses $35,000 plus $3 per unit
Selling price $75 per unit

There was no beginning nor ending inventory. Actual sales were 30,000 units. What is the company's flexible budget operating income (loss) for April?

a) 463,000

b) 550,000

c) 719,000

d) 818,000

e) none of the above
Answer #1

Sales ($75/unit * 30,000) = $2,250,000

Computation of Total Costs

Prime Costs ($35 * 30,000) = ($1,050,000)

Manufacturing Overhead [$137,000 + ($4 * 30,000)] = ($257,000)

Selling & Admin Expenses [$35,000 + ($3 *30,000)] = ($125,000)

Total Costs = ($1,432,000)

Operating Income = $2,250,000 - $1,432,000

Operating Income = $818,000